evaluate the discriminant, and predict the type and number of solutions.
s^2+2s+5=0

Answer by lwsshak3(11628)

evaluate the discriminant, and predict the type and number of solutions.
s^2+2s+5=0
...
a=1, b=2, c=5
discriminant=b^2-4*a*c
=4-4*1*20
=4-20=-16
you will get 2 imaginary solutions